# Comparing Neighbors Together Makes it Easy: Jointly Comparing Multiple Candidates for Efficient and Effective Retrieval

### v1

@@ 4.5 Ablation Study
In Table 4, we explore the impact of different initialization checkpoints for sentence encoder on ZeSHEL reranking performance. We considered three sources for initialization: pre-trained BERT Devlin et al. (2018), ZeSHEL-fine-tuned bi-encoders Yadav et al. (2022), and MS-MARCO-fine-tuned-Cocondenser Guo and Barbosa (2018). For CMC, we find that initializing with a fine-tuned model significantly enhances performance. In contrast, Cocondenser leads to a performance decline in bi-encoders, highlighting CMC’s ability to facilitate domain adaptation of the sentence encoder. In poly-encoder and mixencoder, a similar trend to CMC is found, which can be attributed to their use of a single vector embedding from the candidate. On the other hand, Sum-of-max and Deformer demonstrate fairly consistent performance across various sentence encoders, likely because they use entire token embeddings, making them less sensitive to initialization choices.
@@ 5 Conclusion
In this paper, we present Comparing Multiple Candidates (CMC) which offers a novel approach to retrieve and rerank framework, addressing key issues in scalability and runtime efficiency. By utilizing language models for independent encoding and leveraging the self-attention layer, CMC achieves a balance of speed and effectiveness. Its ability to pre-compute candidate representations offline significantly reduces latency, making it a practical solution for enhancing end-to-end performance. Extensive experimentation validates CMC’s effectiveness, marking it as a promising advancement in the field of neural retrieval and reranking.
Bi-Encoder Initialization ZeSHEL
Methods
Vanilla
BERT
Fine-tuned
w/ ZeSHEL
Fine-tuned
w/ MS MARCO
Valid Test
Deformer $\checkmark$ 65.40 63.58
$\checkmark$ 64.42 62.43
$\checkmark$ 57.01 57.46
Sum-of-max $\checkmark$ 59.57 58.37
$\checkmark$ 58.77 57.65
$\checkmark$ 59.15 58.79
Bi-encoder $\checkmark$ 55.45 52.94
$\checkmark$ 49.32 44.01
Poly-encoder $\checkmark$ 53.37 52.49
$\checkmark$ 55.75 54.22
$\checkmark$ 57.41 58.22
MixEncoder $\checkmark$ 58.63 57.92
$\checkmark$ 58.52 57.68
$\checkmark$ 58.52 57.70
CMC(Ours) $\checkmark$ 56.15 55.34
$\checkmark$ 58.04 56.20
$\checkmark$ 60.05 59.23
Table 4: Comparison of unnormalized accuracy (U.Acc.) on ZeSHEL dataset over different initialization checkpoints (BERT Devlin et al. (2018), Bi-encoder Yadav et al. (2022) fine-tuned for ZeSHEL, and Cocondenser Guo et al. (2020)) fine-tuned for MS MARCO. We denote the best case for each method as bold.
Appendix A Limitations
In the future, we plan to test the CMC’s performance with over 1000 candidates using batch processing. This is because it has not yet been extensively researched whether CMC can effectively retrieve from a large collection, e.g. a collection comprising more than 1 million candidates. Furthermore, we plan to tackle the issue that arises from the concurrent operation of both a bi-encoder and CMC index, which currently requires double the index size. This is a consequence of running two separate encoder models in parallel. To address this, we will investigate various data compression techniques aimed at reducing the space footprint, thereby enhancing the practicality and efficiency of running both the Bi-encoder and CMC simultaneously.
Appendix B Potential Risks
This research examines methods to accelerate the two-stage retrieval and reranking process using efficient and effective CMC. While the proposed CMCmight exhibit specific biases and error patterns, we do not address these biases in this study. It remains uncertain how these biases might affect our predictions, an issue we plan to explore in future research.
Appendix C Detailed Information of Datasets
@@
For standard entity linking, we use AIDA-CoNLL dataset Hoffart et al. (2011) for in-domain evaluation, and WNED-CWEB Guo and Barbosa (2018) and MSNBC Cucerzan (2007) datasets for out-of-domain evaluation. These datasets share the same Wikipedia entity linking set. For comparison with the baseline results from Wu et al. (2020), we employ the 2019 English Wikipedia dump, containing 5.9M entities. We employed a bi-encoder as an initial retriever that yields an average unnormalized accuracy of 77.09 and a recall@10 of 89.21. Unnormalized accuracy is measured for each dataset and macro-averaged for test sets.
Regarding the license for each dataset, AIDA-CoNLL dataset is licensed under a Creative Commons Attribution-ShareAlike 3.0 Unported License. We are not able to find any license information about WNED-CWEB and MSNBC datasets.
@@
ZeSHEL Logeswaran et al. (2019) contains mutually exclusive entity sets between training and test data. The dataset comprises context sentences (queries) each containing a mention linked to a corresponding gold entity description within Wikia knowledge base. The entity domain, also called “world”, varying from 10K to 100K entities, is unique to each domain, testing the model’s ability to generalize to new entities. We employed a bi-encoder from Yadav et al. (2022) whose recall@64 is 87.95. On top of these candidate sets, we report macro-averaged unnormalized accuracy, which is calculated for those mention sets that are successfully retrieved by the retriever and macro-averaged across a set of entity domains. For statistics of entity linking datasets, see Table 5. ZeSHEL is licensed under the Creative Commons Attribution-Share Alike License (CC-BY-SA).
@@
We use a popular passage ranking dataset MS MARCO which consists of 8.8 million web page passages. MS MARCO originates from Bing’s question-answering dataset with pairs of queries and passages, the latter marked as relevant if it includes the answer. Each query is associated with one or more relevant documents, but the dataset does not explicitly denote irrelevant ones, leading to the potential risk of false negatives. For evaluation, models are fine-tuned with approximately 500K training queries, and MRR@10, Recall@1 are used as a metric. To compare our model with other baselines, we employed Anserini’s BM25 Nogueira et al. (2019b). The dataset is licensed under Creative Commons Attribution 4.0 International.
@@
For conversation ranking datasets, we involve The DSTC7 challenge (Track 1) Yoshino et al. (2019) . DSTC 7 involves dialogues taken from Ubuntu chat records, in which one participant seeks technical assistance for diverse Ubuntu-related issues. For these datasets, an official candidate set which includes gold is provided. For details for MS MARCO and DSTC 7 Challenge, see Table 6
@@
Sentence Encoder Intialization BERT-large BERT-base BERT-base BERT-base
@@
Appendix D Training Details
@@
The initial starting point for both the query and candidate encoders can significantly impact performance. The sentence encoders for late interaction models including CMC are initialized using either vanilla huggingface BERT Devlin et al. (2018) or other BERT-based, fine-tuned models. These models include those fine-tuned on the Wikipedia dataset (BLINK-biencoder; Wu et al. (2020)) or MS MARCO (Cocondenser; Gao and Callan (2022)). As cross-encoder is the only model without sentence encoder, we initialize cross-encoder only using vanilla BERT.
We initialize the sentence encoder for CMC and other baselines using (1) vanilla BERT and (2) the BLINK bi-encoder for Wikipedia entity linking datasets, and the MS-MARCO fine-tuned Cocondenser for other datasets. After conducting experiments with both starting points, we selected the best result among them. If more favorable results are found from prior works that conduct reranking over the same candidates, we sourced the numbers from these works.
@@
$\begin{split}\mathcal{L}(q,\tilde{C}_{q})=-\lambda_{1}\sum_{i=1}^{K}y_{i}\log(p_{i})+\lambda_{2}\sum_{i=1}^{K}p_{i}\log\left(\frac{p_{i}}{r_{i}}\right)\\ \end{split}$ (8)
@@
Appendix E Additional Results and Analysis
E.1 Reranking Latency of cross-encoders and CMC
@@
E.2 Effect of Number of Candidates on Retrieval Performance
@@
E.3 Detailed Information of Entity Linking Performance
In Table 9, we present detailed results from Wikipedia entity linking datasets. Also, in table 10, we present detailed results for each world in ZeSHEL test set.
@@ E.3 Detailed Information of Entity Linking Performance
E.4 Ranking Performance on ZeSHEL BM25 candidate sets
@@ Methods
E.5 Ablation Study on Training Strategies
In Table 12, we evaluated the impact of different training strategies on the CMC’s reranking performance on the ZeSHEL test set. The removal of extra skip connections results in only a slight decrease ranging from 0.03 to 0.39 points in normalized accuracy. Also, to examine the effects of a bi-encoder retriever, we remove regularization from the loss. It leads to a performance drop but still shows higher performance than ColBERT, the most powerful baseline in the low latency method. Lastly, we tried to find the influence of negative sampling by using fixed negatives instead of mixed negatives. The result shows a marginal decline in the test set, which might be due to the limited impact of random negatives in training CMC.
@@ E.5 Ablation Study on Training Strategies
E.6 Reranking Performance of Cross-encoders for Various Number of Candidates

### v2

@@ 4.5 Ablation Study
Through the experiments, we notice an improved reranking performance on CMC when transferring the sentence encoder from another domain. To examine whether this is CMC-specific characteristic, we conduct an experiment that investigate how different sentence encoder initializations affect the performance of late-interaction models. For each model, we consider sentence encoder initializations with BERT-based bi-encoders fine-tuned for an in-domain (ZeSHEL; Yadav et al. (2022)) and out-domain (MS-MARCO; Guo and Barbosa (2018)), as well as vanilla BERT Devlin et al. (2018); then for each combination of model and sentence-encoder initialization, we fine-tune the model on ZeSHEL dataset and report its test set results.
In Table 4, different initialization strategies show different effects for each model. CMC and Poly-encoder show significant performance increases with out-of-domain sentence encoder initialization. This can be attributed to both models utilizing single candidate embeddings. Other models, such as Sum-of-max and MixEncoder, show negligible impact from sentence encoder initialization, whereas Deformer and Bi-encoder perform best with vanilla BERT. These findings suggest that CMC and the poly-encoder, which compress candidate information into single embeddings, can benefit from initialization from out-of-domain sentence encoders. As a practical recommendation, we advise practitioners to try out-of-domain initialization when using CMC for potentially improved performance.
(Valid/Test) Sentence Encoder Initialization
Vanilla BERT Fine-tuned with
Model
In-domain
Out-of-domain
(MS MARCO)
Medium- Deformer 65.40/63.58 64.42/62.43 57.01/57.46
Latency Sum-of-max 59.57/58.37 58.77/57.65 59.15/58.79
Low- Bi-encoder 55.54/52.94 55.54/52.94 49.32/44.01
Latency Poly-encoder 53.37/52.49 55.75/54.22 57.41/58.22
MixEncoder 58.63/57.92 58.32/57.68 58.52/57.70
CMC (Ours) 56.15/55.34 58.04/56.20 60.05/59.23
Table 4: Comparison of unnormalized accuracy on valid/test set of ZeSHEL over different sentence encoder initialization (Vanilla BERT Devlin et al. (2018), Bi-encoder fine-tuned for in- Yadav et al. (2022) and out-of-domain Guo et al. (2020)) dataset. We denote the best case for each method as bold.
@@ 5 Conclusion
In this paper, we present a novel and intuitive retrieval and reranking framework, Comparing Multiple Candidates (CMC). By contextualizing the representations of candidates through the self-attention layer, CMC achieves improvements in prediction performance with a marginal increase in speed and memory efficiency. Experimental results show that CMC acts as a seamless intermediate reranker between bi-encoders and cross-encoders. The retrieval pipeline of BE-CMC-CE is not only more accurate but also substantially faster than the widely adopted bi-encoder + cross-encoder (BE-CE). Meanwhile, experiments on four different datasets demonstrate that CMC can serve as the efficient final stage reranker. These empirical results emphasize CMC’s effectiveness, marking it as a promising advancement in the field of neural retrieval and reranking.
Limitations
In the future, we plan to test the CMC’s performance with over 1000 candidates with batch processing. It has not yet been extensively researched whether CMC can effectively retrieve from a large collection, e.g., a collection comprising more than 1 million candidates. Furthermore, we plan to tackle the issue that arises from the concurrent operation of both a bi-encoder and CMC index, which currently requires double the index size. This is a consequence of running two separate encoder models in parallel. To address this, we will investigate an end-to-end training scheme, thereby enhancing the practicality and efficiency of running both the Bi-encoder and CMC simultaneously.
Acknowledgement
We thank Nishant Yadav for his helpful discussions and feedback. This work was supported in part by the National Research Foundation of Korea (NRF) grant (RS-2023-00280883, RS-2023-00222663) and New Faculty Startup Fund from Seoul National University, and with the aid of computing resources from Artificial Intelligence Industry Center Agency, Google cloud platform research credits, and the National Super computing Center with super computing resources including technical support (KSC-2023-CRE-0176).
Appendix A Potential Risks
This research examines methods to accelerate the retrieval and reranking process using efficient and effective CMC. While the proposed CMC might exhibit specific biases and error patterns, we do not address these biases in this study. It remains uncertain how these biases might affect our predictions, an issue we plan to explore in future research.
Appendix B Detailed Information of Datasets
@@
For standard entity linking, we use AIDA-CoNLL dataset Hoffart et al. (2011) for in-domain evaluation, and WNED-CWEB Guo and Barbosa (2018) and MSNBC Cucerzan (2007) datasets for out-of-domain evaluation. These datasets share the same Wikipedia knowledge base. For comparison with the baseline results from Wu et al. (2020), we employ the 2019 English Wikipedia dump, containing 5.9M entities. We employed a bi-encoder as an initial retriever that yields an average unnormalized accuracy of 77.09 and a recall@10 of 89.21. Unnormalized accuracy is measured for each dataset and macro-averaged for test sets.
AIDA-CoNLL dataset is licensed under a Creative Commons Attribution-ShareAlike 3.0 Unported License. We are not able to find any license information about WNED-CWEB and MSNBC datasets.
@@
ZeSHEL Logeswaran et al. (2019) contains mutually exclusive entity sets between training and test data. The dataset comprises context sentences (queries) each containing a mention linked to a corresponding gold entity description within Wikia knowledge base. Unlike Wikipedia entity linking datasets where the entity set of train and test set overlaps, the entity set for ZeSHEL is mutually exclusive and this setup tests the model’s ability to generalize to new entities. We employed a bi-encoder from Yadav et al. (2022) whose recall@64 is 87.95. On top of these candidate sets, we report macro-averaged unnormalized accuracy, which is calculated for those mention sets that are successfully retrieved by the retriever and macro-averaged across a set of entity domains. For statistics of entity linking datasets, see Table 5. ZeSHEL is licensed under the Creative Commons Attribution-Share Alike License (CC-BY-SA).
@@
We use a popular passage ranking dataset MS MARCO which consists of 8.8 million web page passages. MS MARCO originates from Bing’s question-answering dataset with pairs of queries and passages, the latter marked as relevant if it includes the answer. Each query is associated with one or more relevant documents, but the dataset does not explicitly denote irrelevant ones, leading to the potential risk of false negatives. For evaluation, models are fine-tuned with approximately 500K training queries, and MRR@10, Recall@1 are used as a metric. To compare our model with other baselines, we employed Anserini’s BM25 as a retriever Nogueira et al. (2019b). The dataset is licensed under Creative Commons Attribution 4.0 International.
@@
For conversation ranking datasets, we involve The DSTC7 challenge (Track 1) Yoshino et al. (2019) . DSTC 7 involves dialogues taken from Ubuntu chat records, in which one participant seeks technical assistance for diverse Ubuntu-related issues. For these datasets, an official candidate set which includes gold is provided. For statistics for MS MARCO and DSTC 7 Challenge, see Table 6
@@
Appendix C Training Details
@@
The initial starting point for both the query and candidate encoders can significantly impact performance. The sentence encoders for late interaction models including CMC are initialized using either vanilla huggingface BERT Devlin et al. (2018) or other BERT-based, fine-tuned models. These models include those fine-tuned on the Wikipedia dataset (BLINK-bi-encoder; Wu et al. (2020)) or MS MARCO (Cocondenser; Gao and Callan (2022)). As the cross-encoder is the only model without sentence encoder, we initialize cross-encoder using pre-trained BERT (BLINK-cross-encoder; Wu et al. (2020)) or vanilla BERT.
We initialize the sentence encoder for CMC and other baselines using (1) vanilla BERT and (2) the BLINK bi-encoder for Wikipedia entity linking datasets, and the MS-MARCO fine-tuned Cocondenser for other datasets. After conducting experiments with both starting points, we selected the best result among them. If more favorable results for baselines are found from prior works that conduct reranking over the same candidates, we sourced the numbers from these works.
@@
$\begin{split}\mathcal{L}(q,\tilde{C}_{q})&=-\lambda_{1}\sum_{i=1}^{K}y_{i}\log(p_{i})\\ &+\lambda_{2}\sum_{i=1}^{K}p_{i}\log\left(\frac{p_{i}}{r_{i}}\right)\end{split}$ (8)
@@
Appendix D Additional Results and Analysis
D.1 Reranking Latency of cross-encoders and CMC
@@
D.2 Effect of Number of Candidates on Retrieval Performance
@@
D.3 Detailed Information of Entity Linking Performance
In Table 9, we present detailed results for each dataset in Wikipedia entity linking task. Also, in table 10, we present detailed results for each world in ZeSHEL test set.
@@ D.3 Detailed Information of Entity Linking Performance
D.4 Ranking Performance on ZeSHEL BM25 candidate sets
@@ Methods
D.5 Ablation Study on Training Strategies
In Table 12, we evaluated the impact of different training strategies on the CMC’s reranking performance on the ZeSHEL test set. The removal of extra skip connections results in only a slight decrease ranging from 0.03 to 0.39 points in normalized accuracy. Also, to examine the effects of a bi-encoder retriever, we remove regularization from the loss. It leads to a performance drop but still shows higher performance than sum-of-max, the most powerful baseline in the low latency method. Lastly, we tried to find the influence of negative sampling by using fixed negatives instead of mixed negatives. The result shows a marginal decline in the test set, which might be due to the limited impact of random negatives in training CMC.
D.6 Reranking Performance of Cross-encoders for Various Number of Candidates